A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 200-cm and a standard deviation of 0.7-cm. For shipment, 18 steel rods are bundled together.


Find the probability that the average length of a randomly selected bundle of steel rods is between 199.7-cm and 200-cm.

P(199.7-cm < ¯xx¯ < 200-cm) =
Answer #1

Solution:

We are given

µ = 200

σ = 0.7

n = 18

We have to find P(199.7 < x̄ < 200)

P(199.7 < x̄ < 200) = P(x̄<200) - P(x̄<199.7)

Find P(x̄<200)

Z = (x̄ - µ)/[σ/sqrt(n)]

Z = (200 - 200)/[0.7/sqrt(18)]

Z = 0

P(Z<0) = P(x̄<200) = 0.5000

(by using z-table)

Now find P(x̄<199.7 )

Z = (x̄ - µ)/[σ/sqrt(n)]

Z = (199.7 - 200)/(0.7/sqrt(18))

Z = -1.81827

P(Z<-1.81827) = P(x̄<199.7) = 0.034511

(by using z-table)

P(199.7 < x̄ < 200) = P(x̄<200) - P(x̄<199.7)

P(199.7 < x̄ < 200) = 0.5000 - 0.034511

P(199.7 < x̄ < 200) = 0.465489

Required probability = 0.465489